# Efficient Ensemble Conditional Independence Test Framework for Causal Discovery

### v1

@@ Abstract
Constraint-based causal discovery relies on numerous conditional independence tests (CITs), but its practical applicability is severely constrained by the prohibitive computational cost, especially as CITs themselves have high time complexity with respect to the sample size. To address this key bottleneck, we introduce the Ensemble Conditional Independence Test (E-CIT), a general and plug-and-play framework. E-CIT operates on an intuitive divide-and-aggregate strategy: it partitions the data into subsets, applies a given base CIT independently to each subset, and aggregates the resulting p-values using a novel method grounded in the properties of stable distributions. This framework reduces the computational complexity of a base CIT to linear in the sample size when the subset size is fixed. Moreover, our tailored p-value combination method offers theoretical consistency guarantees under mild conditions on the subtests. Experimental results demonstrate that E-CIT not only significantly reduces the computational burden of CITs and causal discovery but also achieves competitive performance. Notably, it exhibits an improvement in complex testing scenarios, particularly on real-world datasets.
@@ 1 Introduction
However, the heavy reliance of constraint-based causal discovery on numerous CITs creates a severe computational bottleneck, significantly limiting practical use. While many studies (Akbari et al., 2021; Mokhtarian et al., 2021; 2023; 2025; Shiragur et al., 2024; Rohekar et al., 2021) have focused on reducing the number of CITs, a more fundamental challenge lies in the high time complexity of CITs themselves (Zhang et al., 2011; Scetbon et al., 2022). Despite some research (Strobl et al., 2019; Schacht & Huang, 2025) on mitigating the cubic time complexity of popular kernel-based conditional independence test (KCIT) (Zhang et al., 2011), Shah & Peters (2018) demonstrate that no single CIT is uniformly effective across all conditional dependence structures. Thus, a critical open question is how to generally reduce the computational cost of CITs while preserving their testing power.
To address this challenge, we propose the Ensemble Conditional Independence Test (E-CIT), a generic plug-and-play framework that can be seamlessly applied to existing CIT methods to mitigate computational burden while maintaining competitive performance. E-CIT adopts an intuitive divide-and-aggregate strategy: given a CIT method, it partitions the dataset into multiple subsets, conducts independent tests on each subset, and aggregates the resulting p-values. For this combination, we introduce a novel method based on the properties of stable distributions, which is theoretically consistent under mild conditions on the subtests, and ensures the reliability of the overall procedure. When the subset size is fixed, this strategy controls the computational complexity of the base CIT to linear in the sample size. The main contributions of this paper are summarized as follows:
  • We introduce E-CIT, a general divide-and-aggregate framework that systematically mitigates the computational complexity of CITs, thereby addressing a fundamental computational bottleneck in causal discovery with respect to sample size.
@@ 2.1 Conditional Independence Testing
We briefly review several representative and recent approaches to CIT, while referring to Li & Fan (2020) for a comprehensive overview. A typical approach in CIT is to define criteria for conditional independence. One of the most widely used measures is Conditional Mutual Information (CMI) (Runge, 2018; Mukherjee et al., 2020; Jamshidi et al., 2024), along with several other metrics (Yu et al., 2020; Wang et al., 2015; Cai et al., 2022). Sen et al. (2017) reformulate CIT as a binary classification problem and apply modern classifiers for hypothesis testing. Recently, the Conditional Randomization Test (CRT) (Candes et al., 2018) has inspired several new methods. For example, Bellot & van der Schaar (2019); Shi et al. (2021) utilize GANs for conditional sampling, whereas Li et al. (2023a; b) employ nearest-neighbor sampling techniques. These methods are particularly effective for handling large conditioning sets.
KCIT (Zhang et al., 2011) is a widely used CIT method leveraging reproducing kernel Hilbert spaces (RKHS), and has inspired many kernel-based extensions (Doran et al., 2014; Scetbon et al., 2022; Zhang et al., 2022; Pogodin et al., 2024). Efforts to accelerate CITs have primarily focused on approximations of KCIT, notably RCIT (Strobl et al., 2019) and FastKCIT (Schacht & Huang, 2025). RCIT employs random Fourier features for efficient approximation, while FastKCIT partitions the dataset with Gaussian mixture models of the conditioning variable $Z$. Although similar in spirit, FastKCIT is specifically tailored to KCIT rather than serving as a general framework like E-CIT.
@@ 2.2 Combination Test
Mudholkar & George (1979) $\sum_{i=1}^{m}\ln\left[p_{i}/(1-p_{i})\right]$
@@
The problem of combining individual p-values into an overall test has long been central in statistics, with important applications in fields such as genomics. Consider the scenario where the same hypothesis is tested $m$ times, yielding $m$ corresponding p-values $p_{1},\dots,p_{m}$. Classical methods for combining these p-values are summarized in Table 1, where $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution. Heard & Rubin-Delanchy (2018) investigate the conditions under which each classical method for combining p-values is most appropriate.
Recently, the problem of combining multiple p-values has received renewed attention (Vovk & Wang, 2020; Geistkemper, 2024), particularly in high-dimensional settings with dependent tests, as is common in biostatistics. Liu & Xie (2020) proposed a Cauchy-based method using inverse probability weighting, which performs well under such conditions (Long et al., 2023). Building on this technique, Liu et al. (2024) developed an ensemble testing method, and Ling & Rho (2022) further generalized this approach using stable distributions. While conceptually related, our work differs in both application domain and underlying modeling assumptions. Those methods are explicitly designed for Whole-Genome Sequencing (WGS) association studies. In that context, their parametric assumptions, such as that the base test statistics follow a normal distribution (Liu & Xie, 2020; Ling & Rho, 2022), are justified. In contrast, the E-CIT framework is designed as a general-purpose tool for the non-parametric context of CIT, where such parametric assumptions are generally not applicable.
@@ 3.1 Ensemble Conditional Independence Test Framework
Figure 1: The framework of E-CIT. Each scatter plot displays samples of variables $X$ and $Y$, with color indicating the value of $Z$. Despite smaller subset sizes, the dependence (black contours) and conditional independence given $Z$ (blue contours) remain clearly distinguishable.
Given $n$ samples from a joint distribution over $X$, $Y$, and $Z$, our goal is to test whether $X\perp\!\!\!\perp Y\mid Z$ using an arbitrary base CIT method. However, as the sample size $n$ grows, the computational cost of many CIT methods can become prohibitive due to their high complexity. Inspired by ensemble learning, we propose the Ensemble Conditional Independence Test (E-CIT) framework to address this issue, as illustrated in Figure 1. We partition the entire dataset into $K$ subsets of size $n_{k}$, where $n=Kn_{k}$. The base CIT is applied independently to each subset, yielding p-values $\{p_{1},\ldots,p_{K}\}$, which are then combined into a final p-value. When $n_{k}$ is fixed, this ensures that the overall computational cost scales linearly with $n$, regardless of the original complexity of the CIT method.
As shown in Figure 1, appropriately sized subsets can retain sufficient information to capture conditional dependence. However, unlike classical parametric hypothesis tests and their associated p-value combination (Liu & Xie, 2020; Ling & Rho, 2022), CITs have a more complex alternative hypothesis $H_{1}$: conditional dependence. As a result, the distribution of p-values under $H_{1}$ can vary significantly across different data-generating mechanisms and CIT methods. This variability poses a challenge for p-value combination, since the statistical properties of a combination method depend heavily on the alternative distribution of p-values (Heard & Rubin-Delanchy, 2018). Therefore, to ensure broad applicability across scenarios and methods, a flexible aggregation strategy that maintains statistical properties under diverse conditions is essential. In the next section, we propose such a method based on the properties of stable distributions for the E-CIT framework.
@@ 3.2 Combining p-values via Stable Distributions
$\mathbb{E}\left[\exp(iuX)\right]=\begin{cases}\exp\left(-\gamma^{\alpha}|u|^{\alpha}\left[1-i\beta\left(\tan\frac{\pi\alpha}{2}\right)(\operatorname{sign}(u))\right]+i\delta u\right),&\alpha\neq 1\\ \exp\left(-\gamma|u|\left[1+i\beta\frac{2}{\pi}(\operatorname{sign}(u))\log|u|\right]+i\delta u\right),&\alpha=1\end{cases}$
@@
The most important property of stable distributions is their generalization of closure under summation, as the sum of independent stable-distributed random variables remains stable. A detailed illustration of this property is provided in Appendix A. In our method, we utilize the closure property in a specific case, as shown below.
@@
This elegant property gives rise to the name of stable distribution. Building upon this property, we define the core of our proposed method:
@@
Given a set of p-values $p_{1},p_{2},\dots,p_{K}$ from independent and identical subtests $\mathcal{H}$, the ensemble test $\mathcal{H}_{e}\left(\mathcal{H},K;\alpha,\beta,\gamma,\delta\right)$ is defined via the test statistic $T_{e}$ as:
@@
It is important to distinguish our approach from related work such as Ling & Rho (2022). Although both methods utilize the closure property of stable distributions, our method yields a different form of the combined p-value. More importantly, our subsequent theoretical analysis is tailored for the challenges of CITs and, consequently, makes no parametric assumptions on the form of the subtests (such as the normality of subtests’ statistics (Liu & Xie, 2020; Ling & Rho, 2022)). Additionally, existing methods such as Stouffer et al. (1949) can be viewed as special cases of our approach. To formally establish the reliability of our method, we present its key theoretical properties below (see Appendix B for detailed proofs).
@@
The ensemble test $\mathcal{H}_{e}$ (for well-defined subtest p-values) satisfies the following properties:
@@
    Unbiasedness [^1] Unbiasedness in hypothesis testing is defined as the rejection probability under the alternative hypothesis being at least the pre-specified significance level (Lehmann & Romano, 2005), distinct from unbiasedness in estimation.: The ensemble test is unbiased if its subtests are unbiased, meaning the ensemble does not compromise the unbiasedness of the individual subtests.
Theorem 1 ensures that our ensemble test is valid for well-defined base p-values. However, we acknowledge that due to the challenges inherent to CITs, these p-values are often approximate. This implies that the guarantees of Theorem 1 may not hold exactly in practice, a point we discuss further in Appendix F.
@@
Lemma 1 establishes the sufficient condition for the power of our ensemble test to converge to 1. However, the conditions stated are not directly interpretable and may be difficult to verify. Thus, we present the following theorem, which provides more relaxed and practically verifiable conditions for convergence.
@@
Theorem 2 highlights the reliability of the E-CIT framework. It shows that E-CIT not only preserve the consistency of the base CITs but also offer a potential way to improve the power of methods lacking theoretical consistency guarantees. More importantly, while many existing CITs have consistency guarantees (Scetbon et al., 2022; Jamshidi et al., 2024), their underlying assumptions can be difficult to satisfy in complex scenarios (Appendix C). In contrast, the consistency of E-CIT established by Theorem 2 does not directly impose assumptions on the testing scenario itself. Instead, it only requires the individual subtests to be reasonably effective. This property enhances the general applicability of E-CIT in challenging situations (see experimental results in Section 4).
@@
The three conditions in Theorem 2 can be easily satisfied in practice. The first condition requires the performance of individual subtests, specifically that the expected p-value of subtest $\mathcal{H}$ under the alternative hypothesis $H_{1}$ is below the significance level. This condition is mild and can be directly translated to requiring that the power of $\mathcal{H}$ exceeds a threshold determined by $f_{1}$, which can be as low as 0.5 in edge cases. See further illustrations in Appendix D.
The second condition concerns the shape of the p-value distribution under $H_{1}$, requiring the density on the left side of $f_{1}$ to exceed the symmetric value on the right. This is natural since p-values under $H_{1}$ tend to concentrate near 0, and it is automatically satisfied when the first condition holds and p-values are approximated by a Beta distribution (Heard & Rubin-Delanchy, 2018) (Appendix D).
The third condition restricts the stable distribution used in $\mathcal{H}_{e}$. The requirement $\alpha\geq 1$ ensures the tail is no heavier than Cauchy distribution, which aligns with statistical intuition about tail behavior (Liu & Xie, 2020). Although $\beta=\delta=0$ can be relaxed in theory (see Eq.(3), Appendix B.3), we fix them to simplify both the proofs and implementation, while using $\alpha$ to control tail heaviness in practice.
@@
Lastly, we discuss the rationale for maintaining flexibility in our E-CIT framework. Since valid p-values follow a uniform distribution under the null hypothesis, the Neyman-Pearson lemma dictates that the uniformly most powerful test statistic for combining p-values should correspond to a monotonic transformation of $\sum_{k=1}^{K}\log f_{1}(p_{k})$ (Casella & Berger, 2024; Heard & Rubin-Delanchy, 2018). Therefore, the ensemble test is optimal when
@@
where $g$ is an arbitrary monotonic function. However, for CIT, unlike traditional parametric tests, the conditional dependence structure of the data under the alternative hypothesis $H_{1}$ can lead to different distributions of $p_{k}^{H_{1}}$ even with the same CIT method. Therefore, our E-CIT allows the flexibility of adjusting $\alpha$ to concisely control $F_{S}$, allowing the test statistic to satisfy the above condition as closely as possible. However, we acknowledge that the theoretically optimal choice of $\alpha$ is context-dependent and requires further analysis for specific CIT methods (see Appendix F for a discussion).

### v2

# arXiv:2509.21021v2[cs.LG] 27 Feb 2026

@@ Abstract
Constraint-based causal discovery relies on numerous conditional independence tests (CITs), but its practical applicability is severely constrained by the prohibitive computational cost, especially as CITs themselves have high time complexity with respect to the sample size. To address this key bottleneck, we introduce the Ensemble Conditional Independence Test (E-CIT), a general-purpose and plug-and-play framework. E-CIT operates on an intuitive divide-and-aggregate strategy: it partitions the data into subsets, applies a given base CIT independently to each subset, and aggregates the resulting p-values using a novel method grounded in the properties of stable distributions. This framework reduces the computational complexity of a base CIT to linear in the sample size when the subset size is fixed. Moreover, our tailored p-value combination method offers theoretical consistency guarantees under mild conditions on the subtests. Experimental results demonstrate that E-CIT not only significantly reduces the computational burden of CITs and causal discovery but also achieves competitive performance. Notably, it exhibits an improvement in complex testing scenarios, particularly on real-world datasets.
@@ 1 Introduction
However, the heavy reliance of constraint-based causal discovery on numerous CITs creates a severe computational bottleneck, significantly limiting its practical use. While many studies (Akbari et al., 2021; Mokhtarian et al., 2021; 2023; 2025; Shiragur et al., 2024; Rohekar et al., 2021) have focused on reducing the number of CITs to streamline the discovery process, a more fundamental challenge lies in the high time complexity of CITs themselves (Zhang et al., 2011; Scetbon et al., 2022). Despite some research (Strobl et al., 2019; Schacht and Huang, 2025) on mitigating the cubic time complexity of the popular kernel-based conditional independence test (KCIT) (Zhang et al., 2011), Shah and Peters (2018) demonstrate that no single CIT is uniformly effective across all conditional dependence structures. Thus, a critical open question is how to generally reduce the computational cost of CITs while preserving their testing power.
To address this challenge, we propose the Ensemble Conditional Independence Test (E-CIT), a general-purpose plug-and-play framework that can be seamlessly applied to existing CIT methods to mitigate the computational burden while maintaining competitive performance. E-CIT adopts an intuitive divide-and-aggregate strategy: given a CIT method, it partitions the dataset into multiple subsets, conducts independent tests on each subset, and aggregates the resulting p-values. For this combination, we introduce a novel method based on the properties of stable distributions, which is theoretically consistent under mild conditions on the subtests, and ensures the reliability of the overall procedure. When the subset size is fixed, this strategy controls the computational complexity of the base CIT to linear in the sample size.
The main contributions of this paper are summarized as follows:
  • We introduce E-CIT, a general-purpose divide-and-aggregate framework that systematically mitigates the computational complexity of CITs, thereby addressing a fundamental computational bottleneck in causal discovery with respect to sample size.
@@ 2.1 Conditional Independence Testing
We briefly review several representative and recent approaches to CIT, while referring to Li and Fan (2020) for a comprehensive overview. A typical approach in CIT is to define criteria for conditional independence. One of the most widely used measures is Conditional Mutual Information (CMI) (Runge, 2018; Mukherjee et al., 2020; Jamshidi et al., 2024), along with several other metrics (Yu et al., 2020; Wang et al., 2015; Cai et al., 2022). Sen et al. (2017) reformulate CIT as a binary classification problem and apply modern classifiers for hypothesis testing. Recently, the Conditional Randomization Test (CRT) (Candes et al., 2018) has inspired several new methods. For example, Bellot and van der Schaar (2019); Shi et al. (2021) utilize GANs for conditional sampling, whereas Li et al. (2023a; b) employ nearest-neighbor sampling techniques. These methods are particularly effective for handling large conditioning sets.
KCIT (Zhang et al., 2011) is a widely used CIT method leveraging reproducing kernel Hilbert spaces (RKHS), and has inspired many kernel-based extensions (Doran et al., 2014; Scetbon et al., 2022; Zhang et al., 2022; Pogodin et al., 2024). Efforts to accelerate CITs have primarily focused on approximations of KCIT, notably RCIT (Strobl et al., 2019) and FastKCIT (Schacht and Huang, 2025). RCIT employs random Fourier features for efficient approximation, while FastKCIT partitions the dataset with Gaussian mixture models of the conditioning variable $Z$. Although similar in spirit, FastKCIT is specifically tailored to KCIT rather than serving as a general framework like E-CIT.
@@ 2.2 Combination Test
Mudholkar and George (1979) $\sum_{i=1}^{m}\ln\left[p_{i}/(1-p_{i})\right]$
@@
The problem of combining individual p-values into an overall test has long been central in statistics, with important applications in fields such as genomics. Consider the scenario where the same hypothesis is tested $m$ times, yielding $m$ corresponding p-values $p_{1},\dots,p_{m}$. Classical methods for combining these p-values are summarized in Table 1, where $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution. Heard and Rubin-Delanchy (2018) investigate the conditions under which each classical method for combining p-values is most appropriate.
Recently, the problem of combining multiple p-values has received renewed attention (Vovk and Wang, 2020; Geistkemper, 2024), particularly in high-dimensional settings with dependent tests, as is common in biostatistics. Liu and Xie (2020) propose a Cauchy-based method using inverse probability weighting, which performs well under such conditions (Long et al., 2023). Building on this technique, Liu et al. (2024) develop an ensemble testing method, and Ling and Rho (2022) further generalize this approach using stable distributions. Notably, these specific methods are explicitly designed for traditional parametric settings, primarily for whole-genome sequencing (WGS) association studies.
@@ 3.1 Ensemble Conditional Independence Test Framework
Figure 1: Overview of the E-CIT framework. Each scatter plot displays samples of variables $X$ and $Y$, with color indicating the value of $Z$. Despite smaller subset sizes, the marginal dependence (black contours) and conditional independence given $Z$ (blue contours) remain clearly distinguishable.
Given $n$ samples from a joint distribution over $X$, $Y$, and $Z$, our goal is to test whether $X\perp\!\!\!\perp Y\mid Z$ using an arbitrary base CIT method. However, as the sample size $n$ grows, the computational cost of many CIT methods can become prohibitive due to their high complexity. Inspired by ensemble learning, we propose the Ensemble Conditional Independence Test (E-CIT) framework to address this issue, as shown in Figure 1. We partition the entire dataset into $K$ subsets of size $n_{k}$, where $n=Kn_{k}$. The base CIT is applied independently to each subset, yielding p-values $\{p_{1},\ldots,p_{K}\}$, which are then combined into a final p-value. When $n_{k}$ is fixed, this ensures that the overall computational cost scales linearly with $n$, regardless of the original complexity of the CIT method.
As illustrated in Figure 1, subsets of sufficient size can effectively capture conditional dependence. However, unlike classical parametric hypothesis tests and their associated p-value combination methods (Liu and Xie, 2020; Ling and Rho, 2022), CITs have a more complex alternative hypothesis, leading to p-value distributions that vary significantly across data-generating mechanisms and base CIT methods. This variability poses a challenge for p-value combination, since the statistical properties of a combination method depend heavily on the alternative distribution of p-values (Heard and Rubin-Delanchy, 2018). Therefore, to ensure broad applicability across scenarios and methods, a flexible aggregation strategy that maintains statistical properties under diverse conditions is essential. In the next section, we propose such a method based on the properties of stable distributions for the E-CIT framework.
@@ 3.2 Combining p-values via Stable Distributions
$\mathbb{E}\left[\exp(iuX)\right]=\begin{cases}\exp\left(-\gamma^{\alpha}|u|^{\alpha}\left[1-i\beta\left(\tan\frac{\pi\alpha}{2}\right)(\operatorname{sign}(u))\right]+i\delta u\right),&\alpha\neq 1\\ \exp\left(-\gamma|u|\left[1+i\beta\frac{2}{\pi}(\operatorname{sign}(u))\log|u|\right]+i\delta u\right),&\alpha=1.\end{cases}$
@@
The most important property of stable distributions is their generalization of closure under summation, as the sum of independent stable-distributed random variables remains stable (detailed in Appendix A). In our method, we utilize the closure property in a specific case, as shown below.
@@
This elegant property gives rise to the name of the stable distribution. Building upon this property, we define the core of our proposed method:
@@
Given a set of i.i.d. p-values $p_{1},p_{2},\dots,p_{K}$ derived from independent and identical subtests $\mathcal{H}$, the ensemble test $\mathcal{H}_{e}\left(\mathcal{H},K;\alpha,\beta,\gamma,\delta\right)$ is defined by the test statistic $T_{e}$:
@@
It is important to distinguish our approach from related work such as Ling and Rho (2022). Structurally, our method is a generalized form of Stouffer et al. (1949), whereas Ling and Rho (2022) generalizes Liu and Xie (2020) and Lipták (1958). More importantly, our subsequent theoretical analysis is tailored for the challenges of CITs and, consequently, makes no parametric assumptions on the form of the subtests (such as the normality of subtests’ statistics (Liu and Xie, 2020; Ling and Rho, 2022)).
To formally establish the reliability of our method, we present its key theoretical properties below (see Appendix B for detailed proofs).
@@
The ensemble test $\mathcal{H}_{e}$ (for exact subtest p-values) satisfies the following properties:
@@
    Unbiasedness [^1] Unbiasedness in hypothesis testing is defined as the rejection probability under the alternative hypothesis being at least the pre-specified significance level (Lehmann and Romano, 2005), distinct from unbiasedness in estimation.: The ensemble test is unbiased if its subtests are unbiased, meaning the ensemble does not compromise the unbiasedness of the individual subtests.
Theorem 1 ensures that our ensemble test is valid for exact base p-values. However, we acknowledge that due to the challenges inherent to CITs, these p-values are often approximate. This implies that the guarantees of Theorem 1 may not hold exactly in practice, a point we discuss further in Appendix F.
@@
Lemma 1 establishes a sufficient condition for the power of our ensemble test to converge to 1. However, the conditions stated are not directly interpretable and may be difficult to verify. Thus, we present the following theorem, which provides more relaxed and practically verifiable conditions for convergence.
@@
Theorem 2 highlights the reliability of the E-CIT framework. It shows that E-CIT not only preserves the consistency of the base CITs but also offers a potential way to improve the power of methods lacking theoretical consistency guarantees. More importantly, while many existing CITs have consistency guarantees (Scetbon et al., 2022; Jamshidi et al., 2024), their underlying assumptions can be difficult to satisfy in complex scenarios (Appendix C). In contrast, the consistency of E-CIT established by Theorem 2 does not directly impose assumptions on the testing scenario itself. Instead, it only requires the individual subtests to be reasonably effective. This property enhances the general applicability of E-CIT in challenging situations (see experimental results in Section 4).
@@
The three conditions in Theorem 2 can be easily satisfied in practice. The first condition requires the performance of individual subtests, specifically that the expected p-value of the subtest $\mathcal{H}$ under the alternative hypothesis $H_{1}$ is below the significance level. This condition is mild and can be directly translated to requiring that the power of $\mathcal{H}$ exceeds a threshold determined by $f_{1}$, which can be as low as 0.5 in edge cases. See further illustrations in Appendix D.
The second condition concerns the shape of the p-value distribution under $H_{1}$, requiring the density on the left side of $f_{1}$ to exceed the symmetric value on the right. This is natural since p-values under $H_{1}$ tend to concentrate near 0, and it is automatically satisfied when the first condition holds and p-values are approximated by a Beta distribution (Heard and Rubin-Delanchy, 2018) (Appendix D).
The third condition restricts the stable distribution used in $\mathcal{H}_{e}$. The requirement $\alpha\geq 1$ ensures the tail is no heavier than Cauchy distribution, which aligns with statistical intuition about tail behavior (Liu and Xie, 2020). Although $\beta=\delta=0$ can be relaxed in theory (see Eq.(3), Appendix B.3), we fix them to simplify both the proofs and implementation, while using $\alpha$ to control tail heaviness in practice.
@@
Lastly, we discuss the rationale for maintaining flexibility in our E-CIT framework. Since valid p-values follow a uniform distribution under the null hypothesis, the Neyman-Pearson lemma dictates that the uniformly most powerful test statistic for combining p-values should correspond to a monotonic transformation of $\sum_{k=1}^{K}\log f_{1}(p_{k})$ (Casella and Berger, 2024; Heard and Rubin-Delanchy, 2018). Therefore, the ensemble test is optimal when
@@
where $g$ is an arbitrary monotonic function. However, for CIT, unlike traditional parametric tests, the conditional dependence structure of the data under the alternative hypothesis $H_{1}$ can lead to different distributions of $p_{k}^{H_{1}}$ even with the same CIT method. Therefore, our E-CIT allows the flexibility of adjusting $\alpha$ to concisely control $F_{S}$, enabling the test statistic to satisfy the above condition as closely as possible. However, we acknowledge that the theoretically optimal choice of $\alpha$ is context-dependent and requires further analysis for specific CIT methods (see Appendix F for a discussion).